# Fibered ribbon pretzels

### v1

@@ 4. Proofs
Suppose there is a pair $e^{j}_{l},e^{j}_{l+1}$ such that neither element appears in $\iota(U)$. Then by simply deleting this pair we obtain an embedding of a new graph $\Gamma^{\prime}$ which differs from $\Gamma_{1}$ only in that the weight of the vertex $v_{j}$ has increased by 2, the weight of $v_{c}$ is either unchanged or increased by 2, and again corresponds to a pretzel knot. As the vertex set is unchanged, yet $\Gamma^{\prime}$ embeds in a lattice of rank rk$(\Gamma)-2$, we arrive at a contradiction of Remark 3.1.
@@
Combining these observations, we see that each $\iota(u^{j}_{1})$ must contain a vector of the form $e^{k}_{*}$ and not $f_{1}$. If this vector is $e^{k}_{1}$ for some $k>1$, then the structure of the $-2$-chain immediately implies that (up to relabelling) $\iota(u^{j}_{i})=e^{k}_{i}-e^{k}_{i+1}$ for each $i$, and so by Observation 4.5, $|n_{k}|=p_{j}$ (recall that all parameters $n$ and $p$ are odd). Moreover, orthogonality implies straight away that $\iota(u^{s}_{1})$ cannot contain $e_{1}^{k}$ for $s\neq j$. It follows that for each $k>1$, there is a unique $U^{j}$ whose embedding contains exactly the vectors in $\iota(v_{k})$. So far we have established that the embedding of the graph $\Gamma_{1}$ is as in Figure 8.
Figure 8. Partial embedding of $\Gamma_{1}$ into the standard negative diagonal lattice with basis $\{e^{1}_{1},\dots,e^{r}_{|n_{r}|},f_{1},f_{2}\}$. The embedding of the remaining $-2$-chains uses only basis vectors of the form $e^{1}_{*}$.
We are left then with $q:=t-(r-1)$ further $U^{j}$’s, the embeddings of which contain only vectors in $\iota(v_{1})$, and such that the embedding of the initial vertex of each contains one of $\{e^{1}_{1},e^{1}_{2},e^{1}_{3}\}$. Appealing again to Observation 4.5, we see that $1\leq q$, and by orthogonality $q\leq 3$. Moreover, as $d+t-r=2$, we see that $q=3-d$; as $d>1$ by assumption, we discard the possibility $q=3$.
@@
It follows that the rank of $\Gamma$ is $|n_{1}|+1$. However, $Q_{\Gamma}$ is diagonally dominant and we have an embedding of $\Gamma$ into a standard negative lattice of rank $|n_{1}|$ wich again contradicts the ‘algebraic version’ of Remark 3.1.
To finish the proof we still need to discard the possibility of an embedding of the lattice $(\Gamma_{2},Q_{\Gamma_{2}})$. This is much easier than the preceding case. In $\Gamma_{2}$ the central vertex has weight $d-r$, which, since $d\geq 1$, yields $|v_{c}\cdot v_{c}|<r$. It is immediate to check, and follows from the arguments above, that we need at least $r$ basis vectors to attach $r$ $-2$-chains of length greater than two. This necessitates a central vertex with $|v_{c}\cdot v_{c}|\geq r$. This contradiction finishes the proof.
Now that we have established Proposition 4.4 we proceed to analyze the case of Type 2B pretzel knots with unitary parameters. Recall from Gabai’s classification, Theorem 2.1, that a Type 2B fibered pretzel knot has an equal number of positive and negative odd parameters. Continuing with our convention of using $d$ for the number of $1$s, $t$ for the number of positive odd parameters and $r$ for the number of negative odd parameters, we can write Gabai’s condition as $d+t=r$. Furthermore, for a Type 2B pretzel knot to be fibered we need either $L^{\prime}=\pm P(2,-2,\dots,2,-2,n)$, $n\in\mathbb{Z}$, which is impossible since in our case $t\neq r$; or $L^{\prime}=\pm P(2,-2,\dots,2,-2,2,-4)$, which is also impossible since this link has an even number of parameters defining it, so $d$ is odd and therefore $t\neq r$, but, at the same time, $L^{\prime}$ has the same number of $-2$ and $2$ entries, which needs $t=r$. We collect this little observation in the next proposition.
@@
As in the Type 2A case, we have two possible negative definite graphs, $\Gamma_{1}$ and $\Gamma_{2}$, associated with these knots. These are depicted in Figure 9 together with the Wu set. The essential difference between Type 3A and Type 2A graphs is that in the former there is one even parameter among the $p_{i}$ and $n_{j}$, while in the latter the unique even parameter had weight ±$\pm 2$ and was considered independently of the other parameters.
@@
The case of $\Gamma_{2}$ is immediate to exclude: there cannot be an embedding of this graph since the linking of the $r$ $-2$-chains, of which there is at most one shortest one of length 2, necessitates a central vertex of weight less or equal than $-r$; however, since $d\geq 1$ we do not have this condition.
@@
Notice that, if the even parameter is negative, the situation described in the second bullet point applies to all $n_{i}\neq n_{i_{0}}$ except for perhaps one exception. Without loss of generality we may assume it applies to all $n_{i}$ with $i\not\in\{i_{0},r\}$; if the even parameter is positive, we may assume it is $p_{1}$ and, just as before, we have at most one $n_{i}$, $i\neq i_{0}$, not falling into the second bullet point. Once again, without loss of generality, we decide this is $n_{r}$. Summing up, except perhaps for the indices $\{i_{0},r\}$, each $n_{i}$ has, through the embedding, a corresponding positive parameter of the same absolute value. There is a rational homology cobordism (see [Lec15, Remark 6.5]) between $Y_{K}([1^{d}],p_{1},\dots,p_{t},n_{1},\dots,n_{r})$ and $Y_{K^{\prime}}([1^{d}],p_{1},\dots,p_{t^{\prime}},n_{i_{0}},n_{r})$, where the positive parameters might have been re-indexed and $t^{\prime}\leq t$. It follows that in order to show that $\Gamma_{1}$ does not embed, it suffices to prove the same statement for $\Gamma^{\prime}_{1}$, which is depicted with partial embedding in Figure 11.
@@
The embedding of $\Gamma^{\prime}_{1}$ implies that $t^{\prime}\leq 4$: there is at most one $-2$-chain of shortest length two and therefore each chain has to be linked to the central vertex with a different basis vector. Moreover, since $K$ and $K^{\prime}$ only differ by a certain number of paired up parameters, for $K^{\prime}$ it holds $0<d=2+2-t^{\prime}$ which forces $t^{\prime}\in\{0,1,2,3\}$. Furthermore, since $K$ and $K^{\prime}$ are cobordant, their signatures coincide, and applying Equation (1) to $K^{\prime}$ and $\Gamma^{\prime}_{1}$ we obtain
@@
  • $t^{\prime}=1$. In this case there is only one $-2$-chain whose embedding will consist solely of basis vectors of the form $e^{r}_{*}$ or $e^{i_{0}}_{*}$. The chain has length $p_{1}-1$ and its embedding necessitates $p_{1}$ basis vectors, while Equation (5) implies that $p_{1}=|n_{i_{0}}|+|n_{r}|-2$. If these vectors were all of the form $e^{r}_{*}$ we have
@@
    and this cannot happen, since we have at least 3 basis vectors in the embedding of $v_{i_{0}}$. Suppose then that they are all of form $e^{i_{0}}_{*}$. By orthogonality $u^{1}_{1}$ has one basis vector in common with $v_{c}$, and similarly $u_{2}$ has either both or neither of its basis vectors in $E:=\{e_{1}^{i_{0}},e^{i_{0}}_{2},e^{i_{0}}_{3}\}$. If neither, then the same is true of each $u_{k}$, $k>1$, so $p_{1}\leq|n_{i_{0}}|-2$. If both, then again the same is true of each $u_{k}$, but then as there are only three basis vectors available the chain has length at most two, and uses only two vectors from $E$. In either case, $|n_{i_{0}}|>p_{1}$, so
@@
  • $t^{\prime}=2$. We now have two $-2$-chains which we need to embed. Using Observation 4.5, we need to use at least one $e^{r}_{*}$, and thus have one $-2$-chain $U^{1}$ embedded with $e^{r}_{*}$ vectors and the other $U^{2}$ with $e^{i_{0}}_{*}$ vectors. By equation (5) we have that there is a total of $|n_{i_{0}}|+|n_{r}|-3$ weight $-2$ vertices and, since there are two chains, we need precisely $|n_{i_{0}}|+|n_{r}|-1$ vectors in their embedding. Since, as remarked in the previous case, $U^{2}$ cannot use all of $U^{i_{0}}_{*}$, it must use all but one of these, while $U^{1}$ uses all of the $e^{r}_{*}$. We may again then invoke the cobordism argument to delete $U^{1}$ and $v_{r}$, and remove $e^{r}_{1}$ from the embedding of $v_{c}$. We are thus left with embedding the graph $\Gamma$ depicted in Figure 12 with the partial embedding shown.
@@
    Figure 12. Partial embedding of graph $\Gamma$ in a lattice of rank $|n_{i_{0}}|$ with one $-2$-chain of length $|n_{i_{0}}|-2$.
@@
    The 3-manifold associated to this graph is a lens space. While it is simple to determine directly, we can appeal to Lisca’s classification [Lis07a, Lemma 7.3] to conclude that $\Gamma$ will embed if and only if $|n_{i_{0}}|=4$ and the length of the $-2$-chain is 2, that is, for the ‘pretzel knot’ $P(1,1,3,-4)$, which is a 2-bridge ribbon knot. For the sake of completeness, in Figure 13 (with $t=0$) we show the ribbon move, yielding two unlinked unknots, on this pretzel knot (with its standard pretzel projection).
@@
    Figure 13. The family of ribbon pretzel knots $P(1,t+1,3,-4-t)$ and a band which describes the ribbon disks. Notice that, when $t=0$, this is a 2-bridge knot.
@@
    Note then that by adding arbitrary paired parameters discarded in the cobordism argument, we arrive at the family $P(1,1,3,-4,q_{1},...,q_{r},-q_{1},...,-q_{r})$, with $q_{i}\geq 3$ odd, $r\geq 0$. That ribbonness is preserved when the paired up parameters are adjacent follows immediately from the cobordism determined by the evident ribbon move (See Figure 1).
@@
  • $t^{\prime}=3$. In this last subcase we have three $-2$-chains and again by Equation (5) a total of $|n_{i_{0}}|+|n_{r}|-3$ vertices of weight $-2$, now needing $|n_{i_{0}}|+|n_{r}|$ basis vectors in their embedding. Arguing as in the preceding point, the $e^{r}_{*}$ must all appear in one $U^{j}$ chain and thus our cobordism argument allows us to again reduce to the case we are dealing with a graph $\Gamma$ just like the one in Figure 12, but with two $-2$-chains attached to the central vertex instead of just one. Just as before, all basis vectors need to be used in the embedding of these two $-2$-chains.
    Again following the arguments of the previous case, we see that each of $u_{1}^{1}$ and $u_{1}^{2}$ has one basis vector in common with those of $v_{c}$ and that, for each $k$, either each vertex in $\{u^{k}_{l}\ |\ l>1\}$ has no basis vectors in common with $v_{c}$, or each vertex has both its basis vectors in common with $v_{c}$. As we must use all of the vectors of $v_{c}$ somewhere, we conclude that one chain satisfies the latter condition, so is length two, with embedding up to relabelling $(e_{1}^{i_{0}}-e_{4}^{i_{0}},-e_{1}^{i_{0}}+e_{3}^{i_{0}})$ (notice the symmetric role played by $e_{1}^{i_{0}}$ and $e_{3}^{i_{0}}$). The other chain, of arbitrary length, uses all the rest of the basis vectors in the embedding of $v_{i_{0}}$, so up to relabelling is $(-e_{2}^{i_{0}}+e_{5}^{i_{0}},-e_{5}^{i_{0}}+e_{6}^{i_{0}},...)$. The knot associated to this graph is now $P(1,t+1,3,-4-t)$ with $t\geq 1$, ribbonness of which again is demonstrated by the same band move as in the previous case (Figure 13). Just as in the preceding point, we may add arbitrary pairs of parameters to arrive at the family $P(1,t+1,3,-4-t,q_{1},\dots,q_{r},-q_{1},\dots,-q_{r})$, with $q_{i}\geq 3$ odd, $r\geq 0,t\geq 1$. Again the obvious ribbon cobordism will reduce the paired parameters when these are adjacent.
@@
Following the straightforward strategy laid out for this article, we now proceed to study the case of fibered Type 3B pretzel knots with unitary parameters which, up to an overall change of sign, we can assume to be positive. If such a knot is fibered, the number of positive and negative parameters coincide. The auxiliary link in Gabai’s Theorem 2.1 has a difference of one between the number of positive and negative entries. It follows that a Type 3B fibered pretzel knot with positive unitary entries has a unique parameter of value 1. Moreover, a cyclic permutation of the rest of the parameters, of which we have an odd number, has them alternating in sign, with the last parameter negative. Some examples of these fibered knots are $P(1,5,-3,-4)$ and $P(1,-3,-7,4,-5,5)=P(1,4,-5,5,-3,-7)$.
@@
Recall we are assuming that our pretzel knots do not have both parameters ±$\pm 1$ and $\mp 2$ since, as explained in Section 2, these knots admit simpler pretzel knot descriptions. This fact will be key in the proof of the next proposition. Once again, since we will be dealing with Seifert spaces in the arguments, for which the order of the parameters in the pretzel knot do not matter, we will consider for Type 3B fibered pretzel knots $K$ the Seifert spaces $Y_{K}(1,p_{1},\dots,p_{t},n_{1},\dots,n_{t+1})$ where $p_{i}\geq 2$, $n_{i}\leq-3$ and there is a unique even parameter.
@@
Type 3B fibered pretzel knots with ±$\pm 1$ parameters are not slice.
@@
We start by establishing the two possible graphs, $\Gamma_{1}$ and $\Gamma_{2}$ associated with $Y_{K}$, the Seifert spaces described in the paragraph preceding this proposition. They are depicted in Figure 14 together with the Wu-set. Notice that the Wu-set and the graphs are independent of which amongst the $p_{i}$ and $n_{j}$ is the even parameter.
@@
The graph $\Gamma_{2}$ does not embed (into a standard negative lattice of the same rank) and therefore the associated pretzel knots are not slice. The lack of embedding comes from the fact that we need a central vertex $v_{c}$ of weight less or equal than $-t-1$ to link the $t+1$ chains of $-2$s of length greater than one, however, $v_{c}\cdot v_{c}=-t$.
@@
  • Each $-2$-chain is linked to the central vertex through a basis vector of the form $e^{i}_{1}$. This linking vector determines that in the embedding of the $-2$-chain only basis vectors of the form $e^{i}_{*}$ appear. Moreover, if a $-2$-chain embeds with $e^{i}_{*}$ vectors, no other $-2$-chain uses these vectors in its embedding. It follows that for each $j=1,\dots,t$, there is a unique $i\in\{1,\dots,t+1\}$ such that $p_{j}\leq|n_{i}|$.
@@
If one follows carefully the proof of the last propostion, one might be left to wonder what happens if we consider the family of Type 3B fibered pretzel knots with ±$\pm 1$ parameters and include the case of $\mp 2$ being the only even parameter. There are indeed many slice, in fact ribbon, fibered pretzels of this sort, one being for example $K=\pm P(1,-3,3,-2)$. The whole family is the obvious one at this point, with an arbitrary number of paired up parameters beyond the 1 and the $-2$. We are discarding these examples since they all admit a simpler presentation as Type 2B slice fibered pretzel knots. In the example proposed we would get $K=\pm P(-3,3,2)$.

### v2

@@ 4. Proofs
[…]
@@
Combining these observations, we see that each $\iota(u^{j}_{1})$ must contain a vector of the form $e^{k}_{*}$ and not $f_{1}$. If this vector is $e^{k}_{1}$ for some $k>1$, then the structure of the 2-2-chain immediately implies that (up to relabelling) $\iota(u^{j}_{i})=e^{k}_{i}-e^{k}_{i+1}$ for each $i$, and so by Observation 4.5, $|n_{k}|=p_{j}$ (recall that all parameters $n$ and $p$ are odd). Moreover, orthogonality implies straight away that $\iota(u^{s}_{1})$ cannot contain $e_{1}^{k}$ for $s\neq j$. It follows that for each $k>1$, there is a unique $U^{j}$ whose embedding contains exactly the vectors in $\iota(v_{k})$. So far we have established that the embedding of the graph $\Gamma_{1}$ is as in Figure 4.
[…]
Figure 4. Partial embedding of $\Gamma_{1}$ into the standard negative diagonal lattice with basis $\{e^{1}_{1},\dots,e^{r}_{|n_{r}|},f_{1},f_{2}\}$. The embedding of the remaining 2-2-chains uses only basis vectors of the form $e^{1}_{*}$.
We are left then with q:=t(r1)q:=t-(r-1) further $U^{j}$’s, the embeddings of which contain only vectors in $\iota(v_{1})$, and such that the embedding of the initial vertex of each contains one of $\{e^{1}_{1},e^{1}_{2},e^{1}_{3}\}$. Appealing again to Observation 4.5, we see that $1\leq q$, and by orthogonality $q\leq 3$. Moreover, as d+tr=2d+t-r=2, we see that q=3dq=3-d; as $d\geq 1$ by assumption, we discard the possibility $q=3$.
@@
It follows that the rank of $\Gamma$ is $|n_{1}|+1$. However, $Q_{\Gamma}$ is diagonally dominant and we have an embedding of $\Gamma$ into a standard negative lattice of rank $|n_{1}|$ which again contradicts the ‘algebraic version’ of Remark 3.1.
To finish the proof we still need to discard the possibility of an embedding of the lattice $(\Gamma_{2},Q_{\Gamma_{2}})$. This is much easier than the preceding case. In $\Gamma_{2}$ the central vertex has weight drd-r, which, since $d\geq 1$, yields $|v_{c}\cdot v_{c}|<r$. It is immediate to check, and follows from the arguments above, that we need at least $r$ basis vectors to attach $r$ 2-2-chains of length greater than two. This necessitates a central vertex with $|v_{c}\cdot v_{c}|\geq r$. This contradiction finishes the proof.
Now that we have established Proposition 4.4 we proceed to analyze the case of Type 2B pretzel knots with unitary parameters. Recall from Gabai’s classification, Theorem 2.1, that a Type 2B fibered pretzel knot has an equal number of positive and negative odd parameters. Continuing with our convention of using $d$ for the number of $1$s, $t$ for the number of odd parameters greater than one and $r$ for the number of negative odd parameters, we can write Gabai’s condition as $d+t=r$. Furthermore, for a Type 2B pretzel knot to be fibered we need either $L^{\prime}=\pm P(2,-2,\dots,2,-2,n)$, $n\in\mathbb{Z}$, which is impossible since in our case $t\neq r$; or $L^{\prime}=\pm P(2,-2,\dots,2,-2,2,-4)$, which is also impossible since this link has an even number of parameters defining it, so $d$ is odd and therefore $t\neq r$, but, at the same time, $L^{\prime}$ has the same number of 2-2 and $2$ entries, which forces $t=r$. We collect this observation in the next proposition.
@@
As in the Type 2A case, we have two possible negative definite graphs, $\Gamma_{1}$ and $\Gamma_{2}$, associated with these knots. These are depicted in Figure 5 together with the Wu set. The essential difference between Type 3A and Type 2A graphs is that in the former there is one even parameter among the $p_{i}$ and $n_{j}$, while in the latter the unique even parameter had weight $\pm 2$ and was considered independently of the other parameters.
[…]
@@
The case of $\Gamma_{2}$ is immediate to exclude: there cannot be an embedding of this graph since the linking of the $r$ 2-2-chains, of which there is at most one shortest one of length 1, necessitates a central vertex of weight less or equal than r-r; however, since $d\geq 1$ we do not have this condition.
@@
Notice that, if the even parameter is negative, the situation described in the second bullet point applies to all $n_{i}\neq n_{i_{0}}$ except for perhaps one exception. Without loss of generality we may assume it applies to all $n_{i}$ with $i\not\in\{i_{0},r\}$; if the even parameter is positive, we may assume it is $p_{1}$ and, just as before, we have at most one $n_{i}$, $i\neq i_{0}$, not falling into the second bullet point. Once again, without loss of generality, we decide this is $n_{r}$. Summing up, except perhaps for the indices $\{i_{0},r\}$, each $n_{i}$ has, through the embedding, a corresponding positive parameter of the same absolute value. There is a rational homology cobordism (see [pretzel, Remark 6.5]) between $Y_{K}([1^{d}],p_{1},\dots,p_{t},n_{1},\dots,n_{r})$ and $Y_{K^{\prime}}([1^{d}],p_{1},\dots,p_{t^{\prime}},n_{i_{0}},n_{r})$, where the positive parameters might have been re-indexed and $t^{\prime}\leq t$. It follows that in order to show that $\Gamma_{1}$ does not embed, it suffices to prove the same statement for $\Gamma^{\prime}_{1}$, which is depicted with partial embedding in Figure 7.
[…]
@@
The embedding of $\Gamma^{\prime}_{1}$ implies that $t^{\prime}\leq 4$: there is at most one 2-2-chain of shortest length one and therefore each chain has to be linked to the central vertex with a different basis vector. Moreover, since $K$ and $K^{\prime}$ only differ by a certain number of paired up parameters, for $K^{\prime}$ we must have $0<d=2+2-t^{\prime}$ which forces $t^{\prime}\in\{0,1,2,3\}$. Furthermore, since $K$ and $K^{\prime}$ are cobordant up to mutation, their signatures coincide, and applying Equation (1) to $K^{\prime}$ and $\Gamma^{\prime}_{1}$ we obtain
@@
  • $t^{\prime}=1$. In this case there is only one 2-2-chain whose embedding will consist solely of basis vectors of the form $e^{r}_{*}$ or $e^{i_{0}}_{*}$. The chain has length $p_{1}-1$ and its embedding necessitates $p_{1}$ basis vectors, while Equation (5) implies that $p_{1}=|n_{i_{0}}|+|n_{r}|-2$. If these vectors were all of the form $e^{r}_{*}$ we have
@@
    and this cannot happen, since we have at least 3 basis vectors in the embedding of $v_{i_{0}}$. Suppose then that they are all of form $e^{i_{0}}_{*}$. By orthogonality $u^{1}_{1}$ has one basis vector in common with $v_{c}$, and similarly $u^{1}_{2}$ has either both or neither of its basis vectors in $E:=\{e_{1}^{i_{0}},e^{i_{0}}_{2},e^{i_{0}}_{3}\}$. If neither, then the same is true of each $u^{1}_{k}$, $k>1$, so $p_{1}\leq|n_{i_{0}}|-2$. If both, then again the same is true of each $u_{k}$, but then as there are only three basis vectors available the chain has length at most two, and uses only two vectors from $E$. In either case, $|n_{i_{0}}|>p_{1}$, so
@@
  • $t^{\prime}=2$. We now have two 2-2-chains which we need to embed. Using Observation 4.5, we need to use at least one $e^{r}_{*}$, and thus have one 2-2-chain $U^{1}$ embedded with $e^{r}_{*}$ vectors and the other $U^{2}$ with $e^{i_{0}}_{*}$ vectors. By Equation (5) we have that there is a total of $|n_{i_{0}}|+|n_{r}|-3$ weight 2-2 vertices and, since there are two chains, we need precisely $|n_{i_{0}}|+|n_{r}|-1$ vectors in their embedding. Since, as remarked in the previous case, $U^{2}$ cannot use all of $U^{i_{0}}_{*}$, it must use all but one of these, while $U^{1}$ uses all of the $e^{r}_{*}$. We may again then invoke the cobordism argument to delete $U^{1}$ and $v_{r}$, and remove $e^{r}_{1}$ from the embedding of $v_{c}$. We are thus left with embedding the graph $\Gamma$ depicted in Figure 8 with the partial embedding shown.
[…]
@@
    Figure 8. Partial embedding of graph $\Gamma$ in a lattice of rank $|n_{i_{0}}|$ with one 2-2-chain of length $|n_{i_{0}}|-2$.
@@
    The 3-manifold associated to this graph is a lens space. While it is simple to determine directly, we can appeal to Lisca’s classification [lisca, Lemma 7.3] to conclude that $\Gamma$ will embed if and only if $|n_{i_{0}}|=4$ and the length of the 2-2-chain is 2, that is, for the ‘pretzel knot’ P(1,1,3,4)P(1,1,3,-4), which is a 2-bridge ribbon knot. For the sake of completeness, in Figure 9 (with $t=0$) we show the ribbon move, yielding two unlinked unknots, on this pretzel knot (with its standard pretzel projection).
[…]
@@
    Figure 9. The family of ribbon pretzel knots P(1,t+1,3,4t)P(1,t+1,3,-4-t) and a band which describes the ribbon disks. Notice that, when $t=0$, this is a 2-bridge knot.
@@
    Note then that by adding arbitrary paired parameters discarded in the cobordism argument, we obtain $P(1,1,3,-4,q_{1},...,q_{r},-q_{1},...,-q_{r})$, with $q_{i}\geq 3$ odd, $r\geq 0$. That ribbonness is preserved when the paired up parameters are adjacent follows immediately from the cobordism determined by the evident ribbon move (See Figure 1).
@@
  • $t^{\prime}=3$. In this last subcase we have three 2-2-chains and again by Equation (5) a total of $|n_{i_{0}}|+|n_{r}|-3$ vertices of weight 2-2, now needing $|n_{i_{0}}|+|n_{r}|$ basis vectors in their embedding. Arguing as in the preceding point, the $e^{r}_{*}$ must all appear in one $U^{j}$ chain and thus our cobordism argument allows us to again reduce to the case we are dealing with a graph $\Gamma$ just like the one in Figure 8, but with two 2-2-chains attached to the central vertex instead of just one. Just as before, all basis vectors need to be used in the embedding of these two 2-2-chains.
[…]
@@
Following the strategy laid out for this article, we now proceed to study the case of fibered Type 3B pretzel knots with unitary parameters which, up to an overall change of sign, we can assume to be positive. If such a knot is fibered, the number of positive and negative parameters coincide. The auxiliary link in Gabai’s Theorem 2.1 has a difference of one between the number of positive and negative entries. It follows that a Type 3B fibered pretzel knot with positive unitary entries has a unique parameter of value 1. Moreover, a cyclic permutation of the rest of the parameters, of which we have an odd number, has them alternating in sign, with the last parameter negative. Some examples of these fibered knots are P(1,5,3,4)P(1,5,-3,-4) and P(1,3,7,4,5,5)=P(1,4,5,5,3,7)P(1,-3,-7,4,-5,5)=P(1,4,-5,5,-3,-7).
@@
Recall we are assuming that our pretzel knots do not have both parameters $\pm 1$ and $\mp 2$ since, as explained in Section 2, these knots admit simpler pretzel knot descriptions. This fact will be key in the proof of the next proposition. Once again, since we will be dealing with Seifert spaces in the arguments, for which the order of the parameters in the pretzel knot do not matter, we will consider for Type 3B fibered pretzel knots $K$ the Seifert spaces $Y_{K}(1,p_{1},\dots,p_{t},n_{1},\dots,n_{t+1})$ where $p_{i}\geq 2$, $n_{i}\leq-3$ and there is a unique even parameter.
@@
Type 3B fibered pretzel knots with $\pm 1$ parameters are not slice.
@@
We start by establishing the two possible graphs, $\Gamma_{1}$ and $\Gamma_{2}$ associated with $Y_{K}$, the Seifert spaces described in the paragraph preceding this proposition. They are depicted in Figure 10 together with the Wu set. Notice that the Wu set and the graphs are independent of which amongst the $p_{i}$ and $n_{j}$ is the even parameter.
[…]
@@
The graph $\Gamma_{2}$ does not embed (into a standard negative lattice of the same rank) and therefore the associated pretzel knots are not slice. The lack of embedding comes from the fact that we need a central vertex $v_{c}$ of weight less or equal than t1-t-1 to link the $t+1$ chains of 2-2s of length greater than one, however, $v_{c}\cdot v_{c}=-t$.
@@
  • Each 2-2-chain is linked to the central vertex through a basis vector of the form $e^{i}_{1}$. This linking vector determines that in the embedding of the 2-2-chain only basis vectors of the form $e^{i}_{*}$ appear. Moreover, if a 2-2-chain embeds with $e^{i}_{*}$ vectors, no other 2-2-chain uses these vectors in its embedding. It follows that for each $j=1,\dots,t$, there is a unique $i\in\{1,\dots,t+1\}$ such that $p_{j}\leq|n_{i}|$.
@@
If one follows carefully the proof of the last proposition, one might be left to wonder what happens if we consider the family of Type 3B fibered pretzel knots with $\pm 1$ parameters and include the case of $\mp 2$ being the only even parameter. There are indeed many slice, in fact ribbon, fibered pretzels of this sort, one being for example $K=\pm P(1,-3,3,-2)$. The whole family is the obvious one at this point, with an arbitrary number of paired up parameters beyond the 1 and the 2-2. We are discarding these examples since they all admit a simpler presentation as Type 2B slice fibered pretzel knots. In the example proposed we would get $K=\pm P(-3,3,2)$.